OpenStudy (anonymous):

Need an answer, to check it... My work included... -5y - 15 = y + 1... (The base question) -5y - 15 = y + 1 - y -y ------------------ -6 -6 -15=1 -15 -15 ------------- -6y = -14 Answer... -6y = -14

OpenStudy (anonymous):

u can't subtract y on both side cause both variables aren't on the same side.

OpenStudy (anonymous):

Omg... Wow. Dangit.... lol Can u teach me how to do it? starting with "-5y - 15 = y + 1"

OpenStudy (anonymous):

Ill try hold on

OpenStudy (anonymous):

Thx and Joel, thx in advance.

OpenStudy (anonymous):

-5y - 15 = y + 1 -5y- y = 15 + 1 -6y = 16 y = 16/-6 = -8/3

OpenStudy (anonymous):

wait joel why would 15 turn into y ?? u gotta show the steps dude

OpenStudy (anonymous):

Yeah I was about to ask that, but I do like the steps... but I just need that one detail

OpenStudy (anonymous):

haha , y doesn't become 15 , it goes to the other side

OpenStudy (anonymous):

oh u re wrote it

OpenStudy (anonymous):

@mikaa_toxica13 , LOL

OpenStudy (anonymous):

So just like magic, poof. Its on the other side? ._. Wow. Math. U r so nice to me.

OpenStudy (anonymous):

hahah , not magic , u need to take it to the other side and u need to change it's sign

OpenStudy (anonymous):

yeaa I did it wrong. sorry can't help

OpenStudy (anonymous):

Hm.. Okay thanks, I got a real question with my test... Ill be sure to double check it and all that... This just so I can learn it... Thanks @mikaa_toxica13 and @JOELMATHEWS1234 ...

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@JOELMATHEWS1234 : Got it ?

OpenStudy (anonymous):

@E.ali , got what ?

OpenStudy (anonymous):

Got any things in question ...:) If u have any problem say :)

OpenStudy (anonymous):

Ummm , @E.ali , I'm unable to understand , what r u trying to say?

OpenStudy (anonymous):

yes if u want :)

OpenStudy (anonymous):

wat s your problem ?